OpenStudy (anonymous):

Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 7 • 6 + 4 = 7 • 10 = 70 B. 10 + (6 ÷ 3) = 10 + 2 = 12 C. 16 – (4 + 7) = 12 + 7 = 19 D. 2 + 5 • 7 = 2 + 35 = 37

Directrix (directrix):

We will test the options.

Directrix (directrix):

7 • 6 + 4 7 • 10 = 70 =42 + 4 = 46 46 is not equal to 70 so this option is NOT correct.

Directrix (directrix):

10 + (6 ÷ 3) = 10 + 2 = 12 @victor.11 You test this one.

OpenStudy (anonymous):

that one is correct

Directrix (directrix):

I agree.

Directrix (directrix):

What about this: 16 – (4 + 7) = 12 + 7 = 19

Directrix (directrix):

@victor.11 ^^^^^^^^^^^^^ Correct or not?

OpenStudy (anonymous):

Use bodmas rule with operands being left associative, b and d are valid

Directrix (directrix):

What is this --> bodmas rule @Nilanjan

OpenStudy (anonymous):

it gives a order how u evaluate a given expression, 'bracket, of ,division , multiplication , subtraction, addition' in short bodmas ,associativity is important as (6/3)/4 is not same as (6/(3/4))